OpenStudy (anonymous):

A 200.0 g copper block absorbs 2.34 × 10^3 J of heat to raise its temperature by 30.0 K. What is the specific heat of copper? Is it 234 something?

OpenStudy (anonymous):

@greenleaf800073

OpenStudy (anonymous):

\[\Delta T\]=30K mass = 200g Q = 2340J C = 2340J/(200g * 30K) C = .39 J/g * K

OpenStudy (anonymous):

T = change in temperature Q = applied energy C = specific heat C = Q/(mass * T)

OpenStudy (anonymous):

awesome

OpenStudy (anonymous):

Heat added = specific heat x mass x (t final - t initial) 2340J = specific heat * 0.2 kg * (50K - 20K) Cu specific heat = 0.39 J /g C

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

A 500.0 g metal block absorbs 3.25 × 103 J of heat to raise its temperature by 50.0 K. What is the substance?

OpenStudy (anonymous):

is this one also asking for specific heat?

OpenStudy (anonymous):

Heat added = specific heat x mass x (t final - t initial) 3250 J = specific heat * 0.5 kg * ( 70K - 20K) specific heat = 0.13 J /g C Looking at a table, Tungsten has 0.134 J /g C , specific heat

OpenStudy (anonymous):

Sorry ..i mean lead, using your chart

OpenStudy (anonymous):

Since it is easy to find the specific heat with the formula above, you then take that number calculated and compare it to the chart you posted

OpenStudy (anonymous):

awesome thanks!! :)

OpenStudy (anonymous):

@roadjester

OpenStudy (anonymous):

its the one about the 500 g metal block

OpenStudy (anonymous):

with the substance table

OpenStudy (roadjester):

what is the variable you're unsure of?

OpenStudy (anonymous):

is it 0.13 J/ g x C or 0.13 j/Kg. K?

OpenStudy (anonymous):

/Kg x K

OpenStudy (anonymous):

he used 0.5 kg so i was curious

OpenStudy (anonymous):

was the original answer correct or was the unit supposed to be written with the kg x K? because on the table there are two ways of writing it at the top

OpenStudy (anonymous):

looks like this might take some time and its 2am where i live so i can just come back in a few hours

OpenStudy (roadjester):

It depends, As long as you are consistent and have a unit of energy divided by a unit of mass and unit of temperature, it's correct.